jigglypuff314 (jigglypuff314):

I'm slightly confused on this one...? \[\large If~~f(x) = \int\limits_{9}^{3x}\sqrt{2t^3+1}dt\]determine the equation of the line tangent to f at x = 3

jigglypuff314 (jigglypuff314):

If you're not too busy would this be "no answer" or something like that? :/ @zepdrix @mathslover @surjithayer help? :3

OpenStudy (anonymous):

That's an interesting problem, and it has a pretty neat answer. What have you tried?

jigglypuff314 (jigglypuff314):

f(3) = intergral from 9 to 9... which already suggests something's weird >,<

jigglypuff314 (jigglypuff314):

truthfully, my first attempt was using second fundamental theorem to get f'(x) then got f'(3) = 114.5906... but still hit the same problem with the f(3) = ?

OpenStudy (anonymous):

I really want to answer but I feel like @zzr0ck3r will ninja me.

OpenStudy (zzr0ck3r):

well now im not sure...

jigglypuff314 (jigglypuff314):

@mathmale could you help me?

OpenStudy (anonymous):

Does that mean I can answer without being ninja'd?

OpenStudy (zzr0ck3r):

\(f'(x)=\sqrt{2(3x)^3+1}\\f'(3)=\sqrt{1459}\\f(3)=0\) so \(0=\sqrt{1459}*3+b \\b=-3\sqrt{1459}\) so \(y=\sqrt{1459}x-3\sqrt{1459}\) is what I think

OpenStudy (anonymous):

^yep.

OpenStudy (anonymous):

's what I got.

jigglypuff314 (jigglypuff314):

>.> so that's why I only got one point off XD ok, so that's whut I answered on my practice test but had forgot f(3) = ? :P mk thanks!

jigglypuff314 (jigglypuff314):

side note @zzr0ck3r forgot chain rule :P 3 sqrt(1459) not just sqrt(1459)

OpenStudy (mathmale):

As you probably realize, \[\large If~~f(x) = \int\limits\limits_{9}^{3x}\sqrt{2t^3+1}dt\] then f(x) is defined as an unfinished definite integral. Anyone know how to find the derivative of the following? \[f(x) = \int\limits\limits_{9}^{x}\sqrt{2t^3+1}dt?\] Look carefully! I made I small change to the definite integral.

OpenStudy (zzr0ck3r):

\(\sqrt{2x^3+1}\) by FTC

OpenStudy (zzr0ck3r):

thats how I got \(f'(x)=\sqrt{2(3x)^3+1}\)

jigglypuff314 (jigglypuff314):

chain rule derivative of 3x = 3 so f'(x) = 3 sqrt (2(3x)^3 + 1)

OpenStudy (zzr0ck3r):

ahh well thats going to change some stuff

zepdrix (zepdrix):

So what's the part you're stuck on? Evaluating f(x) from 9 to 9?

jigglypuff314 (jigglypuff314):

yeah, I had a brain fart :P thought it was und rather than 0

zepdrix (zepdrix):

0 seems right :) You can think of some area under a curve which has no "width" dimension.

OpenStudy (mathmale):

Remember, we're finding the derivative of a function defined as a definite integral. In this particular case, the lower limit of integration, 9, simply disappeasrs. And yes, we must use the chain rule (the derivative of 3x is, indeed, 3). What is the significance/purpose of the result we now have?

OpenStudy (mathmale):

In other words, why do we want the derivative of f(x) in the first place?

jigglypuff314 (jigglypuff314):

equation of line tangent requires slope f'(x) gives slopes of f(x)

OpenStudy (mathmale):

Cool. And, at x=3, that slope is ... ?

jigglypuff314 (jigglypuff314):

3 sqrt(1459) or 114.5906...

OpenStudy (mathmale):

Evaluate \[f'(x)=\sqrt{2(3x)^3+1}\] at x=3. Why? How would we find f(3)?

jigglypuff314 (jigglypuff314):

so equation I got was y = 3 sqrt(1459) (x-3)

OpenStudy (mathmale):

I'm not checking that, but it looks reasonable. Why not leave this as \[3\sqrt{1459}\]without bothering to evaluate it? More accurate to take that approach.

OpenStudy (mathmale):

Good for @zzr0ck3r ! I see that he/she has taken exactly this approach.

jigglypuff314 (jigglypuff314):

I was crunched for time in class while doing it so I left it as what my calculator gave as 114.5906 which was marked correct I understand what I did wrong now, Thank you for all your help! :)

OpenStudy (mathmale):

I thought this was a very good discussion and appreciate all the contributions each of you has made. :)